OpenStudy (anonymous):

Calculus: implicit differentiation/particle's distance changing question? A particle is moving along the parabola 4y = (x + 2)^2 in such a way that its x-coordinate is increasing at a constant rate of 2 units per second. How fast is the particle's distance to the point (-2, 0) changing at the moment that the particle is at the point (2, 4)?

OpenStudy (irishboy123):

i would calculate the velocity vector at (2,4) and dot that with the *unit version* of direction vector from (-2,0) to (2,4). that should give the velocity of the particle along that second vector ie the rate at which the distance is changing between the 2.

OpenStudy (perl):

i would like try out irishboy's solution. the way i would do it is Given: $$\Large { 4y = (x+2)^2\\~\\ 4~\frac{dy}{dt} = 2(x+2)^1 ~ \frac{dx}{dt} \\ ~\\\rm Plug ~in~\frac{dx}{dt}=2 \\ } $$